please answer all these questions and show all working
2. Solve the the following initial value problems. (a) 9y – 12y + 4y = 0) with y(0) = 2 and y(0) = -1 y + 4y + 4y = 0) w
Answer #1

@ gy-124+44 = 0 with yo =2 and y(0)=-1 Characteristic ean gm²-12m+4=0 gm²-6m-6m+4=0 3m (3m- 2-2( 3m-2)=0 (3M-2) (3M-2) = 0from o e² (c-s=2 - arc = 2 e² - 2 e and from e²(-26 + 2 + 2) = ê (-20, +34) = 1 - 2 ct 32= How from $6 20,- 20 = 402 -20 + 3c[d 4y!! + 4y+y = 0 with y(0)=1 & y(0)=2 * 4m ²74m+1=0 am3.2m + 2m +1 -0 am (2m+1)+1(2m+1)=0 (am +1) (2m+1)=0 m= -1 - 2 x .